"Use a quadratic equation to determine the two numbers that satisfy the situation: They differ by 2. The sum of their squares is 130."
Thank you so much, you really help me!

Answer by venugopalramana(3286)

"Use a quadratic equation to determine the two numbers...SAY X AND Y
that satisfy the situation: They differ by 2.
X-Y=2.........................I
The sum of their squares is 130."
X^2+Y^2=130.........................................II
(X+Y)^2+(X-Y)^2=2(X^2+Y^2)
(X+Y)^2+2^2=2*130=260
(X+Y)^2=260-4=256
X+Y=16.....................................III
EQN.I+EQN.III
2X=18
X=18/2=9
Y=16-X=16-9=7
SO X=9 AND Y=7 ARE THE 2 NUMBERS
Thank you so much, you really help me!